OpenStudy (anonymous):

32 pounds = ___ ounces ??

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

16 ounces = 1 pound

OpenStudy (anonymous):

a. 384 b. 8. c. 2 d. 512

OpenStudy (skullpatrol):

Yes, one of them is the correct answer.

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ - guide the asker to a solution.

OpenStudy (skullpatrol):

$$\huge 16 ounces = 1 pound $$

OpenStudy (mathstudent55):

@skullpatrol gave you the information you need. 1 lb = 16 oz 2 * 1 lb = 2 * 16 oz, or 2 lb = 32 oz 3 * 1 lb = 3 * 16 oz, or 3 lb = 48 oz etc. 32 * 1 lb = ... Can you finish it?